# A Review of Scalable and Privacy-Preserving Multi-Agent Frameworks for Distributed Energy Resource Control

### v1

@@ V-B2 Secret sharing-based methods
           (N/A): Not applicable    SD: State decomposition    NI: Noise injection    GC: Garbled circuit    An adversary runs in probabilistic polyn-
           omial time algorithm [250]    Refers to active adversaries
V-C Other miscellaneous and emerging methods
V-C1 State decomposition
V-C2 Noise injection
V-C3 Garbled circuit
VI Future Directions on Scalable and Privacy-Preserving DER Control
With the increasing penetration of DERs, advanced scalable multi-agent control, optimization, and learning frameworks have been developed to adapt to DER-populated power systems. These advancements, driven by the process of DER data across various fields, further increase the power system’s vulnerability to privacy breaches and security concerns. In this section, we extrapolate new approaches for future scalable, privacy-aware, and cybersecure pathways to unlock the full potential of DERs, as well as controlling, optimizing, and learning generic multi-agent cyber-physical systems.
VI-A Improving Accuracy, Privacy, and Algorithm Efficiency
@@
VI-B Establishing Trustworthiness Across Fields
@@
The integration of DERs is creating profound impacts on power grids within the electric power sector, fostering a highly interconnected community with everything as a grid [255]. The highly connected nature of modern power grids requires the transfer of knowledge from different fields to establish strengthened trustworthiness. It requires consideration of coupled cyber-physical power system architecture [256], the interconnected industrial networks [257], and the heterogeneous knowledge from various fields, such as environmental science, human factors and behavioral science, and artificial intelligence (AI).
For example, power grids are transitioning together with the fast-paced progress of AI. The broad spectrum of AI unfolds new possibilities when consolidating power energy resources to achieve greater grid sustainability, resiliency, and security. However, the need to collect, process, and transfer sensitive system and customer data for fine-tuning learning models can raise new technical, economic, and ethical risks that have not been seen before. The privacy and cybersecurity challenges in the AI field are propagating into the electric power sector, e.g., privacy challenges in natural language processing based on machine learning [207] and power system fault diagnosis within quantum computing field [258].
@@
To better manage risks related to individuals, organizations, and society, the U.S. National Institute of Standards and Technology has developed a comprehensive AI risk management framework [259]. Additionally, Majumder et al. [207] point out that privacy and cybersecurity emerge as a paramount concern when integrating large language models (LLMs) into electric energy systems. Besides, emerging PGD-based adversarial attacks can cause devastating attack results for LLMs [260], resulting in catastrophic failures if deployed in power systems. Moreover, by leveraging the principles of quantum mechanics, quantum computing can break widely-used cryptographic systems by making it possible to factor large numbers efficiently. Zhou and Zhang in [261] show the potential of quantum machine learning in providing resilient and secure decision-making of large-scale power systems. The quantum-inspired methods can enhance security via quantum key distribution [262], resist quantum computing attacks, and open new possibilities for data transfer and information processing, e.g., quantum cryptography [263] and quantum communication [264].
Therefore, there is an urgent need to test existing frameworks and develop new privacy-preserving and cybersecurity solutions by incorporating advanced technologies from various fields and domains to benefit the power and energy community. To this end, establishing a standard that can incorporate trustworthiness considerations into the power energy sector, particularly when applying cross-field techniques to DER control problems, presents a meaningful research direction.
VI-C Developing Zero-Trust Standards
@@
VII Conclusion
With the increasing integration of DERs in large-scale power grids, many power system control, optimization, and learning problems require scalable solutions within a multi-agent framework. Besides, the frequent and mandated exchange of sensitive information among agents makes the entire multi-agent system vulnerable to privacy breaches. These privacy breaches can cause privacy and cybersecurity risks to threaten the function of the entire power grid. Therefore, it is crucial to protect privacy and achieve scalability when deploying multi-agent frameworks for DER control, leading to greater sustainability, security, and resilience.
This paper provides a comprehensive review of recent advancements in scalable and privacy-preserving multi-agent frameworks from multi-disciplinary research areas, highlighting their applications for controlling DER in power systems. It offers a systematic summary of multi-agent frameworks based on their scalable computing and information exchange structures, illustrating their applications in DER control problems across different disciplines. This review identifies internal, external, and hierarchical types of adversaries in multi-agent-based DER control problems, including external eavesdroppers, honest-but-curious agents, and system operators and/or coordinators/aggregators. To prevent privacy leakage, this paper further explores mainstream privacy preservation techniques, such as differential privacy, encryption-decryption-based cryptosystem, and Shamir’s secret sharing, along with other and emerging methods such as state decomposition, noise injection, and garbled circuits. Recent advancements underscore the significant scalability and privacy preservation capabilities of these approaches for the electric power sector. Last but not least, this paper discusses three potential research directions on improving accuracy, privacy, and algorithm efficiency, establishing trustworthiness across fields, and developing zero-trust standards.

### v2

@@ 4.3.3 Garbled circuit
Table 5: (Continued)
(N/A): Not applicable    Ref.: Reference    SD: State decomposition    NI: Noise injection    GC: Garbled circuit    An adversary runs in probabilistic polynomial time algorithm [262]    Refers to active adversaries
5 Future Directions on Scalable and Privacy-Preserving DER Control
With the increasing penetration of DERs, advanced scalable multi-agent control, optimization, and learning frameworks have been developed to adapt to DER-rich power systems. These advancements, driven by the process of DER data across various fields, further increase the power system’s vulnerability to privacy breaches and security concerns. In this section, we extrapolate new approaches for future scalable, privacy-aware, and cybersecure pathways to unlock the full potential of DERs, as well as controlling, optimizing, and learning generic multi-agent cyber-physical systems.
5.1 Improving Accuracy, Privacy, and Algorithm Efficiency
@@
5.2 Establishing Trustworthiness Across Fields
@@
The integration of DERs is creating profound impacts on the electric power sector, fostering a highly interconnected community with everything as a grid [263]. The highly connected nature of modern power grids requires the establishment of strengthened ‘trustworthiness’ across different fields. To this end, the definition of trustworthiness across fields can include three key aspects: 1) trustworthy data and features for the artificial intelligence (AI) model; 2) trustworthy analytical results for the DER control; and 3) trustworthy human-machine interaction for the system management.
@@
Trustworthy data and features for the AI model. As power grids transition alongside the rapid progress of AI, the broad capabilities of AI offer new possibilities for consolidating grid-edge DERs to enhance grid sustainability and resiliency. However, the need to collect, process, and transfer sensitive system and customer data for fine-tuning learning models can raise new technical, economic, and ethical risks that have not been seen before. To manage risks, the U.S. National Institute of Standards and Technology has developed a comprehensive AI risk management framework [264] related to individuals, organizations, and society. Therefore, it is essential to securely collect accurate, unbiased, and representative real-world data for AI models to make fair decisions without violating the privacy and security of the system. Trustworthy analytical results for the DER control. The privacy and cybersecurity challenges in the AI field are propagating into the power energy field, e.g., privacy challenges in natural language processing based on machine learning in the electric energy sector [207] and power system fault diagnosis within quantum computing field [265]. Majumder et al. [207] point out that privacy and cybersecurity emerge as a paramount concern when integrating large language models (LLMs) into electric energy systems. Besides, emerging PGD-based adversarial attacks can cause devastating attack results for LLMs [266], resulting in catastrophic failures if deployed in power systems. Moreover, by leveraging the principles of quantum mechanics, quantum computing can break widely-used cryptographic systems by making it possible to factor large numbers efficiently. Zhou and Zhang in [267] show the potential of quantum machine learning in providing resilient and secure decision-making of large-scale power systems. The quantum-inspired methods can enhance security via quantum key distribution [268], resist quantum computing attacks, and open new possibilities for data transfer and information processing, e.g., quantum cryptography [269] and quantum communication [270]. Therefore, there is a need to test existing frameworks and develop new privacy-preserving and cybersecure solutions by incorporating trustworthy analytical results from various fields to benefit the power and energy community. Trustworthy human-machine interaction for the system management. Enhancing the safety and security of human-machine interaction is also critical for making decisions to manage the DER-rich power systems. It requires consideration of coupled cyber-physical power system architecture [271], the interconnected industrial networks [272], and the stochastic human factors on DER control, such as from the behavioral science perspective.
5.3 Developing Zero-Trust Standards
@@
6 Conclusion
With the increasing integration of distributed energy resources (DERs) in large-scale power grids, many power system control, optimization, and learning problems require scalable solutions within a multi-agent framework. Besides, the frequent and mandated exchange of sensitive information among agents makes the entire multi-agent system vulnerable to privacy breaches. These privacy breaches can cause privacy and cybersecurity risks to threaten the function of the entire power grid. Therefore, it is crucial to protect privacy and achieve scalability when deploying multi-agent frameworks for DER control, targeting for greater sustainability, security, and resilience.
This paper provides a comprehensive review of recent advancements in scalable and privacy-preserving multi-agent frameworks from multi-disciplinary research areas, highlighting their applications for controlling DER in power systems. It offers a systematic summary of multi-agent frameworks based on their scalable computing and information exchange structures, illustrating their applications in DER control problems across different disciplines. This review identifies internal, external, and hierarchical types of adversaries in multi-agent-based DER control problems, including external eavesdroppers, honest-but-curious agents, and system operators and/or coordinators/aggregators. Regarding privacy protection, this paper further explores mainstream privacy preservation techniques, such as differential privacy, encryption-decryption-based cryptosystem, and Shamir’s secret sharing, along with other and emerging methods such as state decomposition, noise injection, and garbled circuits. Recent advancements underscore the significant scalability and privacy preservation capabilities of these approaches for the electric power sector. Last but not least, this paper discusses three potential research directions on improving accuracy, privacy, and algorithm efficiency, establishing trustworthiness across fields, and developing zero-trust standards.